# The Dark Side of NFTs: A Large-Scale Empirical Study of Wash Trading

### v2

@@ Abstract.
Blockchain, Ethereum, Non-Fungible Tokens, Wash Trading
@@ 2.1. Ethereum
Based on blockchain technology, Ethereum is a distributed ledger platform with programmable features. With the design of smart contracts and the account-based model, Ethereum allows people to initiate TXNs and develop applications for users to interact with on blockchains (wik, 2023a). Smart contracts (wik, 2023b) are tamper-proof programs stored on the blockchain that run when predetermined conditions are met. They facilitate TXNs in the decentralized system. Ethereum has two account types, i.e., externally-owned account (EOA) and contract account (contributors, 2023). EOAs are controlled by anyone with private keys, while the contract account is associated with the smart contract code (Vujičić et al., 2018). Ether (ETH) is the native token circulating in Ethereum, used as a payment system for verifying TXNs (wik, 2023a). ERC-20 token is a standard for creating alternative tokens developed by smart contracts (OpenZeppelin, 2023). A block TXN (Zheng et al., 2020) constructs the body of the block in Ethereum and refers to the TXN where the sender sends ETH or other tokens to the receiver with some additional information (e.g., smart contract function calls). An internal TXN refers to the TXN that occurs during the execution of a smart contract (Chan and Olmsted, 2017). When a smart contract is called, it may execute multiple functions, and each function may trigger calls to other contracts. The invocation and interaction between these contracts are achieved through internal TXNs. In this paper, we refer to all TXNs involving the transfer of ERC-20 tokens between addresses as ERC-20 token TXNs.
@@ 2.2. NFT Trading
An NFT (Non-Fungible Token) is a unique on-chain identifier that typically refers to digital files. The tokenId and contract address represent a unique NFT on the blockchain. The first transfer event of an NFT is also called the minted event, showing the publishing of an NFT on the blockchain. NFTs are often purchased on various trading platforms, such as OpenSea. Regarding OpenSea, as shown in Figure 1, there are two main ways to purchase an NFT. 1) instant sale, where an NFT is listed at a fixed price, and a buyer can purchase it immediately, with a portion of the trading value being transferred to OpenSea’s official wallet as service fees. 2) auction, where buyers place bids on an NFT, and sellers can accept the offer. In this case, the sellers pay the service fees (part of the auction price) to OpenSea. The payment token for the auction is WETH (Wra, 2023).
@@ 3.1. Type 1: Round-trip Trading
Explanation. Round-trip Trading is the unethical practice of repeatedly buying and selling the same securities to manipulate the finance market (SCOTT, 2023). In the NFT market, Round-trip Trading happens when someone purchases an NFT and promptly resells it, either directly or through multiple addresses, to the original NFT holder. Example. OG:Crystal #4015 was traded 24 times at almost the same price between 0xCF6FF6 and 0xC17D7c within 9 hours, accounting for 126.2 ETH.
@@ 3.2. Type 2: Unprofitable Trading
Explanation. We define that Unprofitable Trading refers to an NFT TXN where the buyer is either funded beforehand by the seller or receives the seller’s return amount, indicating that the TXN is not intended for profit. Normally, the transfer of NFT (seller $\rightarrow$ buyer) and the transfer of TXN amount (buyer $\rightarrow$ seller) occur simultaneously within a block TXN, which is guaranteed by smart contracts. However, Unprofitable Trading occurs with an additional value transfer, e.g., the seller transfers a certain amount of ETH that is similar to the NFT trading price to the buyer shortly before or after the TXN. Another conduction is using a certain amount of ERC-20 tokens with market value instead of ETH. An early value transfer can be interpreted as a form of funding, while the restitution of the amount illustrates that the NFT TXN does not profit from a third party. Example. For Omnimorph #3980, 0x6149ca (seller) transferred 0.23 ETH (104.5% of the trading price) to 0x2beba3 (buyer) three minutes before the NFT sale event. Twenty minutes later, 0x2beba3 (new seller) returned 0.2098 ETH (95% of the trading price) to 0x0f767ef (new buyer) four minutes after the next sale event. The two NFT sellers did not profit from either TXN, which was abnormal for a usual sale event. They conspired to complete the wash trading process.
@@ 3.3. Type 3: Hidden Trading
Explanation. Hidden Trading is third type of wash trading we identify, characterized by collusive actions between sellers and designated buyers. Hidden Trading implies the existence of a series of continuous private tradings where the NFT sellers designate the NFT buyers. On OpenSea, NFT holders can reserve items for specific buyers (Goodman, 2022), i.e., only the addresses approved by the holders have the right to purchase the NFT. Private Tradings disregard the impact of market liquidity and user sentiment because the sellers and buyers know each other before the TXNs occur. Example. Wash traders raised the NFT price from 3.750 ETH, 3.780 ETH to 4 ETH, 7.4 ETH for VeeFriends #7582, all through continuous private tradings.
@@ 3.4. Wash trading pairs/groups
Explanation. We define wash trading pairs as address pairs with high-frequency wash trading behaviors, while a wash trading group consists of multiple relevant wash trading pairs. Suspicious addresses will no longer be regarded as incidentally or deemed innocently involved in NFT wash trading if they frequently participate in Round-trip Trading, Unprofitable Trading, and Hidden Trading. Specifically, address pairs/groups involved in three types of NFT wash trading more than a certain number of times can be marked as wash trading pairs/groups. If there is a common address/TXN connecting both pairs, we assume that the two pairs come from the same wash trading group. Example Loot #2157 was wash-traded ten times between by 0xB7639A and 0x996665 for 1 ETH between 2022-08-10T12:11:14Z and 2022-08-11T03:25:36Z. Shortly after, it was sold by 0xB7639A to 0xcc8990, and the same wash trading behavior happened again between 0xcc8990 and 0xBF1eD4, and more address pairs. Within a day, 178 transactions transpired back and forth at nearly the same price, and the domain names of sellers and buyers were highly similar, all in the form of Chinese license plates. Here, 0xB7639A and 0x996665, 0xcc8990 and 0xBF1eD can be treated as two wash trading pairs, while 0xB7639A, 0x996665, 0xcc8990, and 0xBF1eD is a wash trading group.
@@ 5.1.4. Wash trading confirmation
If a cycle’s D performs in (v1, v2): [e1, e2]; (v2, v3): [e3]; (v3, v1): [e5, e6], the total number of suspicious walks for the cycle is 2 * 1 * 2 = 4. To prevent misjudgment and ensure stringent identification, each cycle is confirmed as Round-trip Trading behavior only when one of the following rules is met. 1) We consider a cycle included in Round-trip Trading when its number of repetitive walks is no less than the strict threshold, 10 * 10 = 1003, meaning at least ten back-and-forth tradings happen for two addresses. or 2) If all the events in one of the walks are sale events, the cycle would be identified as Round-trip Trading.
@@ 5.2. Approach on Unprofitable Trading
ETH and ERC-20 token transfers constitute additional significant evidence for distinguishing wash trading behavior. We establish three rules to detect each Unprofitable Trading behavior. 1) from and to of the related ERC-20 token/block TXNs are the same as the corresponding sale event. 2) Block TXNs with null input data are considered, meaning that they only contain ETH transfers without any invocation of smart contracts. Specifically, the time threshold is 20 minutes3, i.e., only ETH transfers that occurred 20 minutes before/after the sale event will be included. 3) For ERC-20 token transfer, the time threshold is 80 minutes33footnotemark: 3, i.e., only ERC-20 token transfers that occurred 80 minutes before and after the sale event will be included.
@@ 6.1.5. Results of wash trading pairs/groups
6.1.6. Trend of wash trading events.
We investigate the trend for sale/transfer events related to NFT wash trading. Particularly, we exclude OG:Crystal to avoid its extremely leading role in our trend assessment. Figure 4 shows the change in the number of wash trading events of our results from 2021-06-28T15:00:34Z to 2022-06-20:18:28:10Z. The events have grown and become stable since June 2021 and reach their peak on January 11, 2022. It indicates that wash trading has been used consciously after June 2021. For the peak, it is mostly because of the launch of LooksRare (off, 2023), an NFT marketplace with an incentive reward plan that encourages frequent TXNs (LooksRare, 2023) and thus spawns a bunch of wash trading behaviors (Serneels, 2023; La Morgia et al., 2022; Cho et al., 2023) (See 7.1 for details).
@@ 8.1. Threats to validity
Internal validity: 1) Adjusting the support for FP-Growth may affect the experimental results, e.g., the number of wash trading pairs/groups would increase if it is set loosely. Moreover, the support should be reconsidered for different datasets. To ensure the reliability of our results, we set a strict value, i.e., address pairs will be considered suspicious only when they participate in wash trading at least twelve times. 2) Our method may ignore unknown wash trading. We have not yet investigated more complex trading networks, e.g., incorporating internal TXNs to understand the intention of NFT traders from a more underlying perspective and thus reveal more NFT wash trading patterns. However, we are still the most comprehensive study on NFT wash trading.
External validity: The current tools (e.g., OpenSea API) for collecting NFT events have shortcomings, such as mislabeling transfer/sale/minted events. Nevertheless, we address the related issues and make public a guaranteed dataset for researchers.

### v3

@@ Abstract.
Blockchain, Non-Fungible Tokens (NFTs), Wash Trading, Cyber Security
@@ 2. Background
In this section, we introduce the background knowledge of Ethereum and NFTs.
@@ 2.1. Ethereum
Blockchain is a decentralized, peer-to-peer network system that relies on cryptographic algorithms to secure data and consensus mechanisms to validate transactions. These technologies work together to ensure the integrity and transparency of the blockchain, making it a reliable system for various applications (wik, 2023a). Based on blockchain technology, Ethereum is a distributed ledger platform with programmable features. With the design of smart contracts and the account-based model, Ethereum allows people to initiate TXNs and develop applications for users to interact with on blockchains (wik, 2023b). It increases the diversity of the decentralized world, including the prosperity of NFTs.
Smart contracts (wik, 2023c) are tamper-proof programs stored on the blockchain that run when predetermined conditions are met. They facilitate TXNs in the decentralized system. Ethereum has two account types, i.e., externally-owned account (EOA) and contract account (contributors, 2023). EOAs are controlled by anyone with private keys, while the contract account is associated with the smart contract code (Vujičić et al., 2018). Accounts in Ethereum are anonymous but traceable. Moreover, a single user can hold multiple accounts without providing personal information.
Ether (ETH) is the native token circulating in Ethereum, used as a payment system for verifying TXNs (wik, 2023b). ERC-20 token is a standard for creating alternative tokens developed by smart contracts (OpenZeppelin, 2023).
A block TXN (Zheng et al., 2020) constructs the body of the block in Ethereum and refers to the TXN where the sender sends ETH or other tokens to the receiver with some additional information (e.g., smart contract function calls). An internal TXN refers to the TXN that occurs during the execution of a smart contract (Chan and Olmsted, 2017). When a smart contract is called, it may execute multiple functions, and each function may trigger calls to other contracts. The invocation and interaction between these contracts are achieved through internal TXNs. In this paper, we refer to all TXNs involving the transfer of ERC-20 tokens between addresses as ERC-20 token TXNs.
@@ 2.2. NFT Trading
An NFT (Non-Fungible Token) is a unique on-chain identifier that typically refers to digital files, such as photos. Most NFTs are built using the ERC-721 standard (off, 2023a), which defines the primary interface for users to own, trade, and manage their tokens through a smart contract. NFTs are transferred on-chain using functions displayed by Figure 2. The transfer of an NFT involves the from and to addresses, which represent the propagation of the NFT, and the tokenId and contract address, which together represent a unique NFT on the blockchain [^3] We only discuss ERC-721 NFT in this paper.. Within a collection, each tokenId is different. The first transfer of an NFT is also called the minted event, which shows the publishing of an NFT on the blockchain.
NFTs are often purchased on various trading platforms, such as OpenSea. Regarding OpenSea, as shown in Figure 1, there are two main ways to purchase an NFT. 1) instant sale, where an NFT is listed at a fixed price, and a buyer can purchase it immediately, with a portion of the trading value being transferred to OpenSea’s official wallet as service fees. 2) auction, where buyers place bids on an NFT, and sellers can accept the offer. In this case, the sellers pay the service fees (part of the auction price) to OpenSea. The payment token for the auction is WETH (Wra, 2023).
@@ 3. WASH TRADING/TRADER TYPES
In this section, we delve into the high-level ideas of three wash trading types, providing detailed examples. Then, we define wash-trading pairs/groups to flag wash traders and confirm a related case of colluding addresses.
@@ 3.1. Type 1: Round-trip Trading
Explanation. Round-trip Trading is the unethical practice of repeatedly buying and selling the same securities to manipulate the finance market (SCOTT, 2023). In the NFT market, Round-trip Trading happens when someone purchases an NFT and promptly resells it, either directly or through multiple addresses, to the original NFT holder.
Example. OG:Crystal #4015 was traded 24 times at almost the same price between 0xCF6FF6 and 0xC17D7c within 9 hours, accounting for 126.2 ETH.
@@ 3.2. Type 2: Unprofitable Trading
Explanation. We define that Unprofitable Trading refers to an NFT TXN where the buyer is either funded beforehand by the seller or receives the seller’s return amount, indicating that the TXN is not intended for profit. Normally, the transfer of NFT (seller $\rightarrow$ buyer) and the transfer of TXN amount (buyer $\rightarrow$ seller) occur simultaneously within a block TXN, which is guaranteed by smart contracts. However, Unprofitable Trading occurs with an additional value transfer, e.g., the seller transfers a certain amount of ETH that is similar to the NFT trading price to the buyer shortly before or after the TXN. Another conduction is using a certain amount of ERC-20 tokens with market value instead of ETH. An early value transfer can be interpreted as a form of funding, while the restitution of the amount illustrates that the NFT TXN does not profit from a third party.
Example. For Omnimorph #3980, 0x6149ca (seller) transferred 0.23 ETH (104.5% of the trading price) to 0x2beba3 (buyer) three minutes before the NFT sale event. Twenty minutes later, 0x2beba3 (new seller) returned 0.2098 ETH (95% of the trading price) to 0x0f767ef (new buyer) four minutes after the next sale event. The two NFT sellers did not profit from either TXN, which was abnormal for a usual sale event. They conspired to complete the wash trading process.
@@ 3.3. Type 3: Hidden Trading
Explanation. Hidden Trading is third type of wash trading we identify, characterized by collusive actions between sellers and designated buyers. Hidden Trading implies the existence of a series of continuous private tradings where the NFT sellers designate the NFT buyers. On OpenSea, NFT holders can reserve items for specific buyers (Goodman, 2022), i.e., only the addresses approved by the holders have the right to purchase the NFT. Private Tradings disregard the impact of market liquidity and user sentiment because the sellers and buyers know each other before the TXNs occur.
Example. Wash traders raised the NFT price from 3.750 ETH, 3.780 ETH to 4 ETH, 7.4 ETH for VeeFriends #7582, all through continuous private tradings.
@@ 3.4. Wash trading pairs/groups
Explanation. We define wash trading pairs as address pairs with high-frequency wash trading behaviors, while a wash trading group consists of multiple relevant wash trading pairs. Suspicious addresses will no longer be regarded as incidentally or deemed innocently involved in NFT wash trading if they frequently participate in Round-trip Trading, Unprofitable Trading, and Hidden Trading. Specifically, address pairs/groups involved in three types of NFT wash trading more than a certain number of times can be marked as wash trading pairs/groups. If there is a common address/TXN connecting both pairs, we assume that the two pairs come from the same wash trading group.
Example Loot #2157 was wash-traded ten times between by 0xB7639A and 0x996665 for 1 ETH between 2022-08-10T12:11:14Z and 2022-08-11T03:25:36Z. Shortly after, it was sold by 0xB7639A to 0xcc8990, and the same wash trading behavior happened again between 0xcc8990 and 0xBF1eD4, and more address pairs. Within a day, 178 transactions transpired back and forth at nearly the same price, and the domain names of sellers and buyers were highly similar, all in the form of Chinese license plates. Here, 0xB7639A and 0x996665, 0xcc8990 and 0xBF1eD can be treated as two wash trading pairs, while 0xB7639A, 0x996665, 0xcc8990, and 0xBF1eD is a wash trading group.
@@ 4. DATA COLLECTION
In this section, we detail the methodology used to construct our dataset, which is derived from API access and open-source datasets.
@@ 5.1.4. Wash trading confirmation
If a cycle’s D performs in (v1, v2): [e1, e2]; (v2, v3): [e3]; (v3, v1): [e5, e6], the total number of suspicious walks for the cycle is 2 * 1 * 2 = 4. To prevent misjudgment and ensure stringent identification, each cycle is confirmed as Round-trip Trading behavior only when one of the following rules is met. 1) We consider a cycle included in Round-trip Trading when its number of repetitive walks is no less than the strict threshold, 10 * 10 = 1004, meaning at least ten back-and-forth tradings happen for two addresses. or 2) If all the events in one of the walks are sale events, the cycle would be identified as Round-trip Trading.
@@ 5.2. Approach on Unprofitable Trading
ETH and ERC-20 token transfers constitute additional significant evidence for distinguishing wash trading behavior. We establish three rules to detect each Unprofitable Trading behavior. 1) from and to of the related ERC-20 token/block TXNs are the same as the corresponding sale event. 2) Block TXNs with null input data are considered, meaning that they only contain ETH transfers without any invocation of smart contracts. Specifically, the time threshold is 20 minutes4, i.e., only ETH transfers that occurred 20 minutes before/after the sale event will be included. 3) For ERC-20 token transfer, the time threshold is 80 minutes33footnotemark: 3, i.e., only ERC-20 token transfers that occurred 80 minutes before and after the sale event will be included.
@@ 6.1.5. Results of wash trading pairs/groups
6.1.6. Trend of wash trading events.
We investigate the trend for sale/transfer events related to NFT wash trading. Particularly, we exclude OG:Crystal to avoid its extremely leading role in our trend assessment. Figure 5 shows the change in the number of wash trading events of our results from 2021-06-28T15:00:34Z to 2022-06-20:18:28:10Z. The events have grown and become stable since June 2021 and reach their peak on January 11, 2022. It indicates that wash trading has been used consciously after June 2021. For the peak, it is mostly because of the launch of LooksRare (off, 2023b), an NFT marketplace with an incentive reward plan that encourages frequent TXNs (LooksRare, 2023) and thus spawns a bunch of wash trading behaviors (Serneels, 2023; La Morgia et al., 2022; Cho et al., 2023) (See 7.1 for details).
@@ 8.1. Threats to validity
8.1.1. Internal validity
1) Adjusting the support for FP-Growth may affect the experimental results, e.g., the number of wash trading pairs/groups would increase if it is set loosely. Moreover, the support should be reconsidered for different datasets. To ensure the reliability of our results, we set a strict value, i.e., address pairs will be considered suspicious only when they participate in wash trading at least twelve times. 2) Our method may ignore unknown wash trading. We have not yet investigated more complex trading networks, e.g., incorporating internal TXNs to understand the intention of NFT traders from a more underlying perspective and thus reveal more NFT wash trading patterns. However, we are still the most comprehensive study on NFT wash trading.
8.1.2. External validity
The current tools (e.g., OpenSea API) for collecting NFT events have shortcomings, such as mislabeling transfer/sale/minted events. Nevertheless, we address the related issues and make public a guaranteed dataset for researchers. [^5] We have open-sourced our data, both before and after processing in https://drive.google.com/drive/folders/1bddfHZgk3BSmDUN0aTAub7mJ-_1_36ff, which includes transfer and sale events for 285 NFTs collections.
@@ 10. Conclusion and Future Works
In the future, we will develop a real-time visualization interface to track on-chain data, making it available for researchers. This tool aims to provide a more detailed and immediate understanding of NFT transactions and wash trading activities, further supporting research in this area. Additionally, we will collect more known instances of NFT wash trading behavior and design detection algorithms to cover a wider range of such activities. This will enhance our ability to identify and understand the full scope of wash trading in the NFT market.